15. If x is real and: px2 + 2px - 3 < 0 then the possible values of p are ?


 16.if x is real the set of values of x2 + 4x - 2 = 0 is ?



1
Expert's answer
2021-03-01T16:58:31-0500

1).

"px^2 +2px-3\lt0"

if x is real then, solution of x is:

"x=\frac{-2p\pm\sqrt{(2p)^2-4*p*(-3)}}{2p}=\frac{-2p\pm\sqrt{4p^2+12p}}{2p}"


from the above equation, for the real solution of x, there are some restriction to be follow as:

"p\ne0" and

value of p lies in set as: "4p^2+12p\ge0"

"\implies4p^2\ge-12p"

"\implies p\ge-3"


thus, the possible values of p are"[-3, \u221e)-(0)".


2).


"x^2+4x-2=0"

if x is real then the set of values of equation "x^2+4x-2" is:


"x=\frac{-4\pm\sqrt{4^2-4*1*(-2)}}{2*1}=\frac{-4\pm\sqrt{16+8}}{2}=-2\pm\sqrt6"


thus, the set of values of x2 + 4x - 2 = 0 are "[( -2-\sqrt6),(-2+\sqrt6)]"